14. Using the following table, calculate % Response, Cumulative % Response, % Captured Response, Cumulative Captured Response, Lift and Cumulative Lift by Bin/Percentile.

Bin

Percentile

Number of Responders

Number of Non-responders

Number of Observations

1

5

95

280

375

2

10

14

361

375

3

15

12

363

375

4

20

9

366

375

5

25

9

366

375

6

30

8

367

375

7

35

8

367

375

8

40

6

369

375

9

45

5

370

375

10

50

5

370

375

11

55

3

372

375

12

60

1

374

375

13

65

0

375

375

14

70

1

374

375

15

75

1

374

375

16

80

0

375

375

17

85

1

374

375

18

90

0

375

375

19

95

0

375

375

20

100

0

375

375
Answer #1

1. % Respons=R/(U+IS=R)

=12/278+12

=0.0041 %

2.cummulative responses=95+109+121+130+139+147+155+161+166+171+174+175+175+176+177+177+178+178+178+178

=3160